The National Institute of Standards and Technology has built a clock that is off by only 3.3 µs a year. Consider a current of 0.88 A in a wire. how many electrons pass through a cross-section of the wire in 3.3 x 10^-6 s?
1
Expert's answer
2011-02-07T07:25:08-0500
Firstly, we have to determine the charge which is passed in 3.3x10-6 s:
Q = It = 0.88 [A]*3.3 x 10-6 [s] = 2.904 x 10-6 C.
The charge of the electron is e = −1.602176487(40)×10−19
C, then
N = Q/|e| = 2.904 x 10-6 C/1.602176487(40)×10−19
C = 1.813 ×1013
electrons.